We are using a Moving Average Forecast with n=4 and know that F10 = 49. D 10 = 49, and F11 = 53. Find D6 h
Answer #1

F10 = 49

(D6+D7+D8+D9)/4 = 49

D6+D7+D8+D9 = 196 --------------(1)

F11 = 53

(D7+D8+D9+D10)/4 = 53

D7+D8+D9+D10 = 212

Given D10 = 49

D7+D8+D9 = 212 - 49 = 163 ----------------------(2)

From (1) and (2),

D6 = 196 - 163 = 33

D6 = 33

PLEASE LEAVE A LIKE. IT REALLY HELPS ME A LOT. THANK YOU!!!